OpenStudy (anonymous):

Write a indirect proof of the statement, "The lines x+2y=14 and x+2y=-22 are parallel."

OpenStudy (anonymous):

Could I say, "The sum of the triangles sides is 180 degrees.?

OpenStudy (anonymous):

Personally I consider an indirect proof a proof by contradiction. You try to prove the opposite of the statement and then show that its impossible so the original statement must be true So lets start off with assuming that "The lines x+2y=14 and x+2y=-22 are not parallel and the 2 lines intersect." Next we demonstrate that the lines do not intersect whatsoever We set both equations equal to each other and see at what point on the coordinate system do these 2 lines intersect But before we do that we need to set the equation as an x= ... so that both sides equal x #1: x=14-2y #2: x=-22-2y Solve for y: 14-2y=-22-2y 36=0y 36=0 This is absurd and in this case the 2 lines don't intersect. This is a contradiction Hence we conclude that both lines are parallel

OpenStudy (anonymous):

Parallel lines never intersect. If 2 lines are not parallel then they must intersect